OpenStudy (anonymous):

A train travels along a track and its speed (in miles per hour) is given by s(t)=29t for the first half hour of travel. Its speed is constant and equal to s(t)=29/2 after the first half hour. Here time t is measured in hours. How far (in miles) does the train travel in the first hour of travel? I keep getting 14.5 and I have tried multiple ways.

OpenStudy (ddcamp):

What class is this for? Calculus?

OpenStudy (anonymous):

Pre-calculus

OpenStudy (ddcamp):

Hmm... That makes things a bit more difficult. What section are you on?

OpenStudy (anonymous):

our chapter title is called, Integrals, antiderivatives, and the Fundamental Theorem of CalculusIntegrals, antiderivatives, and the Fundamental Theorem of Calculus

OpenStudy (ddcamp):

Ahh, antiderivatives. When working with position, speed, and acceleration, we know that position is the antiderivative of speed. What's the antiderivative of the speed function?

OpenStudy (anonymous):

I am looking at my notes and I do not quite understand the anti derivative so my guess would be 14.5t \[\int\limits_{0}^{1/2}\]?

OpenStudy (ddcamp):

The really simple rule with antiderivatives is: \[ax^n \rightarrow \frac{a}{n+1}x^{n+1}\]

OpenStudy (ddcamp):

So: \[29t^1 \rightarrow \frac{ 29 }{ 2 }t^2 = 14.5t^2\]

OpenStudy (ddcamp):

Usually there would be a +C, but we're assuming it starts at a position of zero.

OpenStudy (anonymous):

so what would you do next?

OpenStudy (ddcamp):

So, we now have a position function that uses time as the input. How far would the train travel in the first 30 minutes?

OpenStudy (anonymous):

so it would be 1/2 right?

OpenStudy (ddcamp):

Yeah, t=1/2

OpenStudy (anonymous):

So then I would solve it right?

OpenStudy (anonymous):

Ah finally got the answer thanks!